OpenStudy (anonymous):

Wendy earns $6.00 an hour babysitting. She babysat 4 hours Friday, 5 hours Saturday, and more hours Sunday. She earned $78.00 for the weekend. How many hours did she baby sit on Sunday? A. 4 B. 5 C. 6 D. 7

OpenStudy (anonymous):

How many hours did she babysit on friday and saturday?

OpenStudy (anonymous):

idk........answer???????

OpenStudy (anonymous):

No. I don't give answers, dammit! I teach people HOW TO THINK! The answer to MY question is 1st grade math. Tell me how many hours did she babysit on friday and saturday?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

No what?

OpenStudy (anonymous):

answer please

OpenStudy (anonymous):

Screw you.

OpenStudy (anonymous):

what

OpenStudy (mertsj):

if she earned 78 dollars in all and got 6 dollars an hour, how many hours did she babysit in all?

OpenStudy (anonymous):

The answer is 4. 6x4- 24 6x5= 30. 78-30-24= 24. 24/6= 4